OpenStudy (anonymous):

The lengths of the sides of a triangle are 8cm,12cm and 6cm. If the shortest side of a similar triangle is 12cm, what is the length of its longest side?

OpenStudy (anonymous):

24 cm 6/12 = 12/x x = 24

OpenStudy (anonymous):

24 cm